at a college baseball game 4/5 of the spectators are home team fans. the rest of the spectators are opposing team fans. there are 750 opposing team fans.
A. find the fraction of spectators who are opposing team fans.
B.Find the total number of spectators at the game.
C.one third of the home team fans at the game attend the college how many home team fans attend the college.

Found 2 solutions by rfer, palanisamy:
Answer by rfer(16322)

A)
1/5
B)
750/.2=3750
C)
3000*1/3=3000/3=1000

Answer by palanisamy(496)

At a college baseball game 4/5 of the spectators are home team fans.
The rest of the spectators are opposing team fans.
A) The fraction of spectators who are opposing team fans = 1-4/5
= (5-4)/5
= 1/5
B) Given 1/5 part of spectators = 750
Therefore Total number of spectators = 750/(1/5)
= 750*5
= 3750
C) Number of home team fans = 3750-750 = 3000
Given, one third of the home team fans at the game attend the college.
Number of home team fans attend the college = 3000*(1/3) = 1000